OpenStudy (anonymous):

What transformation has changed the parent function f(x) = log2x to its new appearance shown in the graph below? logarithmic graph passing through point negative 2, 0. f(x + 3) f(x − 3) f(x) + 3 f(x) − 3 @ganeshie8 @hartnn @mathstudent55

OpenStudy (anonymous):

f(x+3)

OpenStudy (anonymous):

because log2(-2 + 3) = log2(1) = 0

OpenStudy (michele_laino):

please note that: \[f(-2+3)=\log _{2}(-2+3)=\log _{2}(1)=0\]

OpenStudy (michele_laino):

we have got same idea! @M4thM1nd

OpenStudy (anonymous):

Exactly teh same lol

OpenStudy (anonymous):

so A? @M4thM1nd

OpenStudy (anonymous):

Yes

hartnn (hartnn):

please don't give out answers directly! thanks.

OpenStudy (anonymous):

he explained the steps -.- @hartnn

OpenStudy (anonymous):

i explained @hartnn

hartnn (hartnn):

you gave the answer, then explained. thats not how it should work :P

OpenStudy (anonymous):

same thing.. at least ik how it should be done

OpenStudy (anonymous):

considering the time interval between the answer and the explanation, the order does not matter ^^

OpenStudy (anonymous):

Just kidding ;) Merry Christmas everybody

OpenStudy (anonymous):

f(x)=log(x) is a parent function, and log(2x) is a vertical shift up by log(2). (because, log(2x)=log(2)+log(x), ) comparing f(x=log(x) << parent function f(g)=log(x)+log(2) << a very small vertical shift.

OpenStudy (anonymous):

The function f(x) = 20(3)x represents the growth of a fish population every year in a local lake. Jesse wants to manipulate the formula to an equivalent form that calculates every half-year, not every year. Which function is correct for Jesse's purposes? f(x) = 20(3)the x over 2 power f(x) = 20(3 to the one half power)2x f(x) = twenty halves(3)x f(x) = 40(2)x @i

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

I NEED HELP PLZ ANYONE RN

OpenStudy (anonymous):

@cwrw238 @Kainui

OpenStudy (anonymous):

As you want half year, so f(x) = 20*(3)^(x/2) will give the growth of fishes in x half years

OpenStudy (anonymous):

so f(x) = 20*(3)^(x/2)

OpenStudy (anonymous):

@iamabarbiegirl